### v2

# MOOR-0001-2024.00

[…]
MOOR-0001-2024.00
[…]
Aouad, Ji, and Shaposhnik
[…]
The Pandora’s Box Problem with Sequential Inspections
[…]
The Pandora’s Box Problem
with Sequential Inspections
[…]
The Pandora’s box problem (Weitzman, 1979) is a core model in economic theory that captures an agent’s (Pandora’s) search for the best alternative (box). We study an important generalization of the problem where the agent can either fully open boxes for a certain fee to reveal their exact values or partially open them at a reduced cost. This introduces a new tradeoff between information acquisition and cost efficiency. We establish a hardness result and employ an array of techniques in stochastic optimization to provide a comprehensive analysis of this model. This includes (1) the identification of structural properties of the optimal policy that provide insights about optimal decisions; (2) the derivation of problem relaxations and provably near-optimal solutions; (3) the characterization of the optimal policy in special yet non-trivial cases; and (4) an extensive numerical study that compares the performance of various policies, and which provides additional insights about the optimal policy. Throughout, we show that intuitive threshold-based policies that extend the Pandora’s box optimal solution can effectively guide search decisions.
[…]
Pandora’s box problem, dynamic programming, approximate algorithms
1 Introduction
Sequential search problems refer to a class of optimization problems that have long been studied in the economics, operations research (OR), and computer science (CS) literature. In his seminal work, Weitzman, (1979) studies the Pandora’s box sequential search problem,11endnote: 1We refer to this model as the standard Pandora’s problem. in which Pandora is given a collection of boxes containing random prizes drawn from known and independent probability distributions. She can adaptively pay box-dependent fees to open boxes and observe their realizations until she decides to select the prize held within one of the open boxes, or stop exploring without selecting any box. Her objective is to maximize the expected profit, that is, the expected value of the collected prize minus the expected combined box-opening fees. This fundamental model, which captures an agent’s search for the best alternative, found applications in consumer search, project management, job recruitment, and asset-selling problems, to name a few.
In this paper, we consider a generalization of this problem in which Pandora can also exercise a cheaper and possibly less accurate option of partially opening boxes. When partially opening a box, Pandora observes its type, an informative signal based on which she updates her belief about the prize inside the box. We refer to this variant as Pandora’s Box with Sequential Inspection (PSI for short). Similarly to the standard problem, she may only select a prize contained in a box that was fully opened, that is, a box whose exact prize has been observed. However, the ability to partially open boxes introduces a new tradeoff. On the one hand, partially opening a box reveals certain information about the prize value that allows Pandora to potentially detect “bad outcomes” at a lower cost, and thus, avoid paying the full opening fee. On the other hand, while the cost of partial opening is cheaper than the cost of full opening, it may not be negligible, in which case overusing the option to partially open may lead to excessive costs.
To put this model in concrete terms, consider job recruitment, an application discussed in Weitzman, (1979). Recent technological advances transform job recruitment processes in various industries. For example, video conferencing is often used for remote (offsite) interviews. More recently, AI systems have been introduced to conduct such initial remote interviews without human intervention; the transcriptions are then reviewed by human evaluators. Remote interviews provide convenient ways for companies to learn more about candidates at reduced costs, compared with onsite interviews. Nevertheless, remote interviews cannot fully replace onsite visits, which are often still necessary to eventually make job offers. Moreover, while cheaper than onsite visits, substantive remote interviews – beyond the preliminary screenings based on CVs – still require time from human interviewers or evaluators to assess the candidates’ fit for the position. In selecting a candidate for a vacant position, companies must therefore decide how to synthesize the different mediums to identify and select qualified candidates while minimizing costs. Put into the framework of our model, Pandora represents the company, boxes correspond to candidates whose prior evaluations are based on screenings and automated assessments, and box opening actions correspond to conducting interviews – offsite or onsite. The goal is to effectively administer the hiring process by deciding which candidates should be interviewed and in which format: cheaper and less accurate offsite interviews, or more expensive and informative onsite visits.
More generally and beyond the particular application, our work tries to study how to integrate different sequential information acquisition methods in the context of stochastic optimization problems. To our knowledge, this is the first work that examines this feature, which is quite common in various practical sequential search problems (a more detailed discussion can be found in Section 1.2 that surveys related work).
Weitzman, (1979) showed that there exist optimal policies to the standard Pandora’s problem that admit an elegant and simple threshold-based structure. A threshold is defined for each box, indicating the order in which boxes should be considered for opening, reducing the problem to an optimal stopping time problem. The decision about when to stop is determined using the same thresholds by comparing the largest prize amongst open boxes to the largest threshold amongst closed boxes; if the incumbent prize is larger, Pandora should stop; otherwise, she should proceed to the next box, and so forth. However, it has been observed in the literature (see Section 1.2) that seemingly small changes in the model assumptions break the structure of the optimal policy and lead to problems whose optimal solution cannot be analytically characterized. Our goal is to study this fundamental sequential search process when partial information acquisition is possible. Specifically, we strive to answer the following questions: How should partial opening and full opening be prioritized? Could threshold-based policies, in the same vein as those defined in the standard Pandora’s box problem, describe the optimal policy in settings with sequential inspections? Beyond the characterization of optimal policies, we are also interested in developing simple and intuitive policies with provably good performance guarantees.
1.1 Results and contributions
We summarize below the main findings and contributions of our work.
Modeling. We introduce a new model for sequential search by an agent, which incorporates sequential information acquisition into stochastic optimization problems. Such information acquisition process can be represented by a directed acyclic graph describing the information states of each box (see Figure 1). Our graph layout is unique in that boxes can be inspected repeatedly and there is more than one way to inspect boxes. To our knowledge, this generalized graph layout has not been previously studied. Moreover, it raises a fundamental technical challenge, which has been a long-lasting open question in the literature (we further discuss this in Section 1.2).
Hardness. We prove that computing optimal policies for PSI is NP-hard, contrary to the standard Pandora’s box problem. The proof is based on a reduction defined using PSI instances that consist of boxes with binary prizes and where partial opening indicates one of two types.
Characterization of an optimal policy. Despite the hardness result, we identify threshold-based rules that almost fully describe the optimal policies. Similarly to the Pandora’s box problem, one type of threshold is associated with decisions to fully open boxes. To capture the added value of partial inspections, we define two additional types of thresholds associated with decisions to partially open boxes. We show that, in most cases, the optimal actions can be prescribed using the newly defined thresholds. In particular, we characterize the optimal stopping criterion (i.e., when to stop searching and collect the best prize presently available), and identify sufficient conditions under which it is optimal to fully or partially open a box. We identify special non-trivial cases where our characterization fully describes an optimal policy, and show that a simple policy, which is defined using these thresholds, is asymptotically optimal. Numerical experiments show that optimal actions are identified by our theoretical characterization in more than 92% of all systems states, on average, and this percentage increases in the number of boxes of the problem instance.
Simple robust policies. Next, we focus on the class of committing policies.22endnote: 2Our approach is inspired by the class of committing algorithms studied by Beyhaghi and Kleinberg, (2019), with the notable difference that our setting involves sequential inspections. These semi-adaptive policies specify a priori whether each box should be partially or fully opened the first time it is inspected. Interestingly, each committing policy induces a corresponding multi-armed bandit problem (MAB) in which the optimal indices coincide with our proposed thresholds. By leveraging fundamental results on the adaptivity gap of stochastic submodular optimization problems, we argue that the optimal a committing policy can be efficiently approximated within a ratio of 11/e1-1/e.
Numerical experiments. We conduct extensive numerical experiments to evaluate the performance of various policies on a diverse set of synthetic instances and obtain additional insights about the optimal policy:
  • Solving the problem to optimality using exact dynamic programming methods is computationally prohibitive, even when the number of boxes is around 10, the prize distributions have a small support, and the number of types is small. These computational challenges warrant the development of simple suboptimal policies.
  • The number of times an optimal policy exercises the partial opening option increases as (i) there are more boxes, (ii) the partial opening thresholds tend to be larger than the corresponding full opening thresholds, and (iii) the distributions of the prizes are similar across boxes. In contrast, when boxes tend to be different, much about the order by which boxes are inspected is determined without partial opening.
  • The newly introduced thresholds are key to the development of effective policies. Specifically, the class of committing policies, which utilize our thresholds, achieves a robust performance. In particular, the best committing policy is very close to the optimum, both in terms of average and worst-case performance. In line with the asymptotic analysis, we observe that an intuitive threshold-based committing policy33endnote: 3Here, the policy commits to partial opening vs. full opening simply by comparing the corresponding thresholds for each box. is near-optimal when the number of boxes is sufficiently large.
Our proofs combine several methods used in recent literature, including sample path coupling and interchange arguments (Attias et al., 2017), reductions to submodular maximization problems (Beyhaghi and Kleinberg, 2019), and results based on the information relaxation duality approach (Balseiro and Brown, 2019). In this context, our main technical contribution lies in analyzing threshold-based rules using coupling arguments. We also reveal an unnoticed connection between approximation methods independently proposed in the literature. Specifically, we show that the cost accounting technique of Beyhaghi and Kleinberg, (2019) amounts to a relaxation of Whittle’s integral, a more general upper bound for multi-armed bandit superprocesses.
Overall, our results suggest that generalized Weitzman’s thresholds play an important role in characterizing both optimal and near-optimal inspection and selection decisions when the decision-maker can sequentially inspect boxes. This aspect of selection processes has not been thoroughly studied in previous literature, as we discuss next.
1.2 Related Work
Our work relates to several streams of literature in economics, operations research, and computer science. We divide them imperfectly into three parts: generalization of the standard problem, Markovian MAB problems, and applications of sequential search.
The Pandora’s box problem and its variants. The seminal work of Weitzman, (1979) led to substantial research on more general and realistic settings. For example, Vishwanath, (1992) consider a generalization with parallel searches, where Pandora can open several boxes simultaneously at a discounted cost. The authors identify distributional assumptions under which it is optimal to open boxes according to Weitzman’s thresholds. Olszewski and Weber, (2015) examine a generalization where the objective function depends on all discovered prizes.44endnote: 4In the standard Pandora’s box problem, the objective function is simply the maximum observed prize. They show that if a threshold-based rule is unconditionally optimal, then the problem is reducible to the standard Pandora’s box setting. Boodaghians et al., (2020) study a variant of the Pandora’s box problem in which precedence constraints are imposed on the inspections of boxes. They show that when the precedence constraints are tree-shaped, the optimal policy can still be described as a threshold-based rule. Gibbard, (2022) extend the Pandora’s box problem by assuming that the prize is additive and that a two-step sequential inspection reveals the exact value of each part. Conceptually, such a model studies how information acquisition should be prioritized when the acquisition process can be broken into multiple steps. The model is a special case of the MAB problem (Bertsimas and Nino-Mora, 1996) that inherits the indexability structure of the optimal policy. In contrast, our model attempts to capture a different tradeoff between two information acquisition methods (a cheaper and less precise versus a more expensive and accurate), a characteristic which we show to be challenging analytically and computationally.
Recently, there has been a growing interest in non-obligatory inspections (Attias et al., 2017, Doval, 2018, Beyhaghi and Kleinberg, 2019) where Pandora may select any closed box without inspecting it (in which case the prize is random at selection). Building on a clever cost accounting method, Beyhaghi and Kleinberg, (2019) show that this setting is a special case of maximizing a stochastic non-negative submodular function under a matroid constraint, a breakthrough that paved the way to the design of approximation algorithms. Our algorithmic results in Section 4 draw from this line of reasoning. Closer to our work, Beyhaghi, (2019) studies a setting in which Pandora can choose from multiple opening actions for each box. However, unlike our work, the search process in that scenario does not feature the notion of sequential inspections. Therefore, such a model does not capture the key tradeoff which we aim to study between reducing uncertainty significantly at a higher cost, and reducing it only partially at a lower cost with an option to further reduce the uncertainty later at an overall higher cost. This very tradeoff is present in the applications that motivate our work. Finally, we note the work of Clarkson et al., (2020) that study a problem where an object is hidden at one of several locations according to some known distribution. The goal is to find it in the minimum expected number of attempts using two modes of inspection, fast or slow. This can be seen as a variant of the Pandora’s box problem, where the reward is binary, and boxes are correlated and can be inspected indefinitely.
Markovian multi-armed bandits. From a technical perspective, the standard Pandora’s box problem is a special case of the Markovian MAB (Gittins, 1979), in which an agent sequentially activates (pulls) one out of many bandits (arms). Arms correspond to independent Markov reward processes, which generate a reward and transition to a new state upon activation. The optimal solution admits an index structure, known as the Gittins index55endnote: 5Gittins indices and thresholds in Weitzman’s policy are equivalent in the special case of the Pandora’s box problem (see, e.g., an early version of Olszewski and Weber, 2015)., where every state of every bandit is associated with an index, and it is always optimal to activate the arm whose current state’s index is maximal. However, the optimality of the Gittins index breaks under small variations of the model assumptions, such as correlated arms (Pandey et al., 2007), switching costs (Asawa and Teneketzis, 1996), or delayed feedback (Eick, 1988). We similarly observe that index policies are suboptimal in our setting (see Example 2.2 in Section 2.2).
The problem considered in our paper can be viewed as a special case of MAB superprocesses (Gittins et al., 2011), in which there exist multiple ways to pull each arm. That is, after an arm is selected, one has to choose the mode of activation, which affects transitions and rewards. General knowledge about this class of problems is rather limited. Unlike the standard MAB problem, MAB superprocesses need not admit an optimal index policy. Much of the existing literature explores special cases where index policies are “unambiguously” optimal (Whittle, 1980), as well as heuristic solution methods and upper bounds (Brown and Smith, 2013, Balseiro and Brown, 2019). The fact that we identify that our problem is NP-Hard suggests that a complete characterization of the optimal policy is not possible, not even through the MAB framework.
Sequential search and applications. The stylized sequential search process introduced by Weitzman, (1979) has been widely used in economics, marketing, and operations research. This includes empirical work (De los Santos et al., 2012, Ursu, 2018), behavioral research (Gabaix et al., 2006), and economic models that use the Pandora’s box as a discrete-choice framework (Miller, 1984, Robert and Stahl, 1993, Choi et al., 2018, Derakhshan et al., 2020, Chu et al., 2020, Immorlica et al., 2020). Our work offers a richer modeling structure to represent decision-making processes where alternatives can be inspected in stages with varying costs, although our current contribution remains theoretical.
We note in passing about the existence of a literature on sequential search problems modeled using a Bayesian framework where there is a single box that can be sampled repeatedly for a fee and the decision maker tries to decide when to stop sampling with the goal of maximizing the expected value of the highest observed sample minus the sampling costs (see, e.g., the recent work of Baucells and Zorc, 2023, and the references therein). Although the idea of sequential inspection is similar to our work, the key characteristics of our model, a tradeoff between cheap and costly inspections across multiple alternatives, are not captured by this literature.
2 Problem Formulation
Pandora (the decision-maker) is given $N\in\mathbb{N}^{*}\textsection$ boxes containing random prizes. Each box $i\in[N]$ is associated with a priori unknown prize $V_{i}$ and type $T_{i}$ drawn from a known joint probability distribution $\mathcal{D}_{i}$. We assume that $V_{i}$ is either a continuous or discrete scalar random variable and that $T_{i}$ is a discrete scalar random variable with support $\Gamma_{i}$.66endnote: 6Throughout, we use capital letters to denote random variables and lowercase letters to denote their realizations. We assume that the support of $\mathcal{D}_{i}$ is contained in $\mathbb{R}^{+}\times\Gamma_{i}$, and that the random variables $(V_{1},T_{1}),\ldots,(V_{N},T_{N})$ are mutually independent (for each box $i$, the random variables $V_{i}$ and $T_{i}$ are dependent; otherwise the type would not be informative about the prize). Pandora can open a box in one of two inspection modes: full opening (F-opening) and partial opening (P-opening). The realization of type $T_{i}$ is revealed immediately by P-opening box $i\in[N]$, while the corresponding F-opening action reveals the realization of $(V_{i},T_{i})$. These opening actions incur constant costs, respectively denoted by $c_{i}^{F}$ and $c_{i}^{P}$, which are also known to Pandora.
Figure 1 illustrates the state evolution of each box following various actions. We call the initial state of a box closed; here, the realizations of $V_{i}$ and $T_{i}$ are unknown. P-opening a closed box changes its state to partially open. In this case, Pandora observes the realization of $T_{i}$, which we denote by $t_{i}\in\Gamma_{i}$, and going forward, she updates her belief about $V_{i}$ to the conditional distribution $\operatorname{Pr}[V_{i}\in\cdot\mid T_{i}=t_{i}]$. We sometimes refer to this conditional distribution as $V_{i}\mid t_{i}$ for short. We say that a box is fully open once F-opening takes place. Here, Pandora observes the realizations of $V_{i}$ and $T_{i}$, which we denote by $v_{i},t_{i}$, respectively. Lastly, the problem terminates when a fully open box is selected. We note that full opening is a prerequisite for selection in a way similar to the standard Pandora’s box problem. The state of the system is jointly described by the collective individual state of all boxes.
An adaptive policy in our generalized Pandora’s box problem is a mapping between system states and feasible actions. There are three types of feasible actions: (1) P-opening a closed box, (2) F-opening a closed or partially open box, and (3) selecting a fully open box. Alternatively, this can be seen as a cascade process of (1) deciding whether to select a box (stop) or keep inspecting (continue), (2) choosing a box to inspect, and (3) deciding on a box-opening mode. Our goal is to develop a policy that maximizes the expected profit, defined as the selected prize minus the box-opening costs.
2.1 Dynamic programming (DP) formulation
The problem naturally lends itself to a DP formulation. While the state space is exponentially large (and therefore, the DP cannot be approached using standard solution methods), we exploit this formulation to derive structural properties of the optimal policy.
State space. Each state of the DP can be described by a tuple $(\mathcal{C},\mathcal{P},y)$, where $\mathcal{C}\subseteq[N]$ is the set of closed boxes, $\mathcal{P}$ is formed by all pairs $(i,t_{i})$, describing each P-opened box $i$ and its corresponding type $t_{i}$ (revealed by P-opening), and $y$ is the maximal prize found so far (it is easy to see that all fully open boxes whose prize is smaller than $y$ can be discarded, since at most one prize can be claimed). We observe that the total number of states is at least $\prod_{i=1}^{N}\left(1+\mid\Gamma_{i}\mid\right)$, which is exponentially large.
Actions. At each state $(\mathcal{C},\mathcal{P},y)$, we can fully open a closed or partially open box, partially open a closed box, or select a fully open box and stop. After F-opening a box $i\in\mathcal{C}$ and observing the realization $v_{i}$ of its prize $V_{i}$, we transition to state $(\mathcal{C}\setminus\{i\},\mathcal{P},\max\{y,v_{i}\})$. Hereinafter, we slightly abuse the notation by using $i\in\mathcal{P}$ in place of $(i,t_{i})\in\mathcal{P}$. If we choose to F-open a box $i\in\mathcal{P}$, we transition to state $(\mathcal{C},\mathcal{P}\setminus\{(i,t_{i})\},\max\{y,v_{i}\})$, where $v_{i}$ is the realization of $V_{i}|t_{i}$ revealed by our action. Likewise, if we choose to P-open a box $i\in\mathcal{C}$, we transition to state $(\mathcal{C}\setminus\{i\},\mathcal{P}\cup\{(i,t_{i})\},y)$, where $t_{i}$ is the type realization of $T_{i}$ revealed by P-opening box $i$. Finally, when we select the fully open box whose value is equal to $y$, the decision process terminates.
Bellman’s Equation. Let $J(\mathcal{C},\mathcal{P},y)$ denote the value function in state $(\mathcal{C},\mathcal{P},y)$ under an optimal policy. Then, $J(\mathcal{C},\mathcal{P},y)$ satisfies the Bellman’s Equation:
$J(\mathcal{C},\mathcal{P},y)$ $=$ $\max\left\{\begin{array}[]{ll}y&\text{select (stop)}\\ -c_{i}^{F}+\mathbb{E}_{V_{i}}\left[J(\mathcal{C}\setminus\{i\},\mathcal{P},\max\{V_{i},y\})\right]&\text{F-open box }i\in\mathcal{C}\\ -c_{i}^{F}+\mathbb{E}_{V_{i}\mid T_{i}=t_{i}}\left[J(\mathcal{C},\mathcal{P}\setminus\{(i,t_{i})\},\max\{V_{i},y\})\right]&\text{F-open box }i\in\mathcal{P}\\ -c_{i}^{P}+\mathbb{E}_{T_{i}}\left[J\left(\mathcal{C}\setminus\{i\},\mathcal{P}\cup\{(i,T_{i})\},y\right)\right]&\text{P-open box }i\in\mathcal{C}\end{array}\right$. (5)
$J(\emptyset,\emptyset,y)$ $=$ $y.$
In what follows, we use the shorthand $s_{0}=([N],\emptyset,0)$ to designate the initial system state.
Policies. A feasible policy $\pi$ is a mapping from each state to a corresponding action, which is selected amongst all those that are feasible in the state in question. For simplicity, we assume that there exists a unique optimal policy that satisfies Bellman’s Equation.
2.2 Illustrative examples
To provide some intuition about why P-opening could be beneficial, consider the following simple examples.
Example 2.1
There is a large number of i.i.d. boxes; each contains a prize of 110 with probability $\frac{1}{2}$ and zero otherwise. The box-opening costs are $c_{i}^{F}=100$ and $c_{i}^{P}=\epsilon$, and P-opening acquires the same information as F-opening (nevertheless, F-opening is still required to collect the prize). It is easy to see that given only the F-opening inspection mode, it is not worthwhile to F-open any box and the expected profit of any policy that opens boxes is negative. However, given the option to P-open, one could identify a box that contains the prize prior to F-opening it. In that case, the expected profit is 110100=10110-100=10 minus $\epsilon$ times the expected number of opened boxes, which is equal to two.
This example shows how the option to P-open could transform a problem where opening boxes is not worthwhile into one in which it is. Hence, the optimality gap for the naive policy that ignores P-opening (i.e., Weitzman’s policy, which is a natural candidate for Pandora’s box-like problems) is unbounded.
Example 2.2
Consider the state $(\{1\},\emptyset,y)$ where there is one box left with two types: ‘exceptional’ and ‘average,’ which can be revealed by P-opening. Box 1 is ‘exceptional’ with probability $\delta=0.01$ and is ‘average’ with probability $1-\delta$. If the box is exceptional, the prize is known to be $2$ almost surely, but if it is average, the prize is uniformly distributed in $[0,1]$. Assuming that the F-opening cost is $c_{1}^{F}=\frac{1}{3}$ and the P-opening cost is $c_{1}^{P}=\frac{1}{3}\delta$, it can be easily seen that when $y=0$, both inspection modes are preferred over stopping (which yields 0). Moreover, F-opening dominates P-opening because F-opening is inevitable in this case. However, when $y=1$, F-opening would be no longer worthwhile compared to stopping, because the cost $c_{1}^{F}=\frac{1}{3}$ exceeds the expected gain $\delta\cdot(2-y)=\delta$. By contrast, P-opening (followed by the F-opening) of an exceptional box is advantageous because the expected cost is $c_{1}^{P}+\delta c_{1}^{F}=\frac{1}{3}\delta+\frac{1}{3}\delta=\frac{2}{3}\delta$, whereas the expected gain is $\delta$.
Intuitively, in the context of job hiring, the instance captures a candidate who is likely ‘average’ but has a small chance of being ‘exceptional.’ A low-cost online interview can reveal this. If no good candidate has been found ($y=0$), it is optimal to proceed directly with an onsite interview. However, if a strong candidate has already been identified ($y=1$), the onsite visit should be reserved for an ‘exceptional’ applicant, identified via a relatively short online screening.
Technically, the example shows that the optimal policy is not indexable. If it were, each action would receive a fixed index, and the best action—corresponding to the highest index—would be independent of $y$, which is not the case in Example 2.2. In MAB terms, the optimal activation of one arm depends on the state of others. This illustrates the added complexity of bandit superprocesses, which do not inherit the indexability of classic MABs (see also Section 1.2).
3 Characterizations of Optimal Policies
We begin by studying structural properties of optimal policies, which can be exploited to identify optimal actions using well-designed opening thresholds. In Section 3.1, we show that it is optimal to stop when the largest available prize exceeds all opening thresholds. In Section 3.2, we establish that whenever the largest opening threshold corresponds to an F-opening action, it is optimal to F-open the corresponding box. In contrast, we show that the complementary case, in which the largest opening threshold corresponds to a P-opening action, is more complex to analyze, as there exists no threshold-based decision rule that is universally optimal. Nevertheless, in Section 3.3, we identify sufficient conditions for which P-opening the box with the largest threshold is optimal. In Section 3.4, we provide a complete characterization of the optimal policy for a special case of a PSI with binary prizes where partial opening reveals two types.
3.1 Opening thresholds and optimal stopping rule
We start by defining opening thresholds, a type of index associated with each potential action. Although there may not exist an optimal index policy (as shown in Section 2.2), we show that such indices explain to a large degree the optimal policy. While applicable to general system states, the thresholds can be most easily illustrated by considering the special case where there is a single box that is not fully open. This setting also allows us to express the optimal stopping criterion using our thresholds. We let $x^{+}=\max\{0,x\}$ for every $x\in\mathbb{R}$.
Definition 3.1 (Thresholds)
For every box $i\in[N]$, we define the opening thresholds:
  • If box $i$ is closed, the F-threshold $\sigma^{F}_{i}$ is a solution to
    $c_{i}^{F}=\mathbb{E}_{V_{i}}\Big[(V_{i}-\sigma^{F}_{i})^{+}\Big].$ (6)
  • If box $i$ is partially opened with type $t_{i}$, the (conditional) F-threshold $\sigma^{F\mid t_{i}}_{i}$ is a solution to
    $c_{i}^{F}=\mathbb{E}_{V_{i}\mid T_{i}=t_{i}}\Big[(V_{i}-\sigma^{F\mid t_{i}}_{i})^{+}\Big].$ (7)
  • If box $i$ is closed, the P-threshold $\sigma^{P}_{i}$ is a solution to
    $c_{i}^{P}=\mathbb{E}_{T_{i}}\Big[\max\{0,-c_{i}^{F}+\mathbb{E}_{V_{i}\mid T_{i}}\big[(V_{i}-\sigma^{P}_{i})^{+}\big]\}\Big].$ (8)
  • If box $i$ is closed, the FP-threshold $\sigma^{F/P}_{i}$ is a solution to
    $c_{i}^{P}=\mathbb{E}_{T_{i}}\Big[\max\{0,c_{i}^{F}-\mathbb{E}_{V_{i}\mid T_{i}}\big[(V_{i}-\sigma^{F/P}_{i})^{+}\big]\}\Big].$ (9)
    Observe that the threshold is well-defined even in the case that $c_{i}^{P}\geq c_{i}^{F}$, in which case P-opening is not beneficial.
Assuming the prize distributions are non-atomic, it is not difficult to see that the above thresholds $\sigma_{i}^{F}$, $\sigma_{i}^{F\mid t_{i}}$, $\sigma_{i}^{P}$, and $\sigma_{i}^{F/P}$ exist and are uniquely defined. Moreover, there is a natural ordering between threshold values:
Observation 1 (Thresholds order)
  1. 1.
    For any box $i$, either $\sigma^{F/P}_{i}=\sigma^{F}_{i}=\sigma^{P}_{i}$, $\sigma^{F/P}_{i}<\sigma^{F}_{i}\leq\sigma^{P}_{i}$, or $\sigma^{P}_{i}\leq\sigma^{F}_{i}<\sigma^{F/P}_{i}$;
  2. 2.
    For any box $i$, we have $\underset{t_{i}\in\Gamma_{i}}{\min}~\sigma_{i}^{F|t_{i}}\leq\sigma_{i}^{F/P}$ and $\underset{t_{i}\in\Gamma_{i}}{\max}~\sigma_{i}^{F|t_{i}}\geq\sigma_{i}^{P}$.
Economic interpretation.
The set of F-thresholds $\{\sigma_{i}^{F}\}$ in Equation (6) describes the optimal index policy for the standard Pandora’s box problem (Weitzman, 1979). The conventional economic interpretation for this opening threshold is that of a fair strike price of a hypothetical option contract. That is, if $c_{i}^{F}$ is the cost of a call option to buy prize $V_{i}$ with strike price $\sigma_{i}^{F}$, then $\sigma_{i}^{F}$ is the “fair” strike price that equalizes the cost and the expected profit of buying this option. This intuition carries over to our problem where inspections are sequential using the notion of a compound option or, informally, a “two-stage call option.” This analogy is developed in Appendix 8.
Single-box setting and optimal stopping rule.
We justify our definition of the opening thresholds through their ability to express optimal policies in single closed-box settings, and to prescribe an optimal stopping rule in general. Namely, the single closed-box setting is the collection of states $({\cal C},{\cal P},y)$ where there is only one closed box that can be inspected, i.e., $\mathcal{C}=\{i\}$ for some box $i\in[N]$ and $\mathcal{P}=\emptyset$. The difference between the expected profit from F-opening and from stopping is exactly
$c_{i}^{F}+\mathbb{E}\Big[\max\left\{V_{i},y\right\}\Big]-y=-c_{i}^{F}+\mathbb{E}\Big[\max\{V_{i}-y,0\}\Big]=-\mathbb{E}\Big[\max\{V_{i}-\sigma_{i}^{F},0\}\Big]+\mathbb{E}\Big[\max\{V_{i}-y,0\}\Big]\ .$ (10)
The F-threshold can be therefore interpreted as a break-even point for the value of $y$ under which F-opening is preferred to stopping, and over which stopping is preferred. Similarly, in this single-box case, the difference between the expected profit from P-opening (and following the optimal action thereafter) and from stopping is
$-c_{i}^{P}+\mathbb{E}_{T_{i}}\Big[\max\{y,-c_{i}^{F}+\mathbb{E}_{V_{i}\mid T_{i}}\big[\max\left\{V_{i},y\right\}\big]\}\Big]-y$ (11)
$=$ $-c_{i}^{P}+\mathbb{E}_{T_{i}}\Big[\max\{0,-c_{i}^{F}+\mathbb{E}_{V_{i}\mid T_{i}}\big[\max\left\{V_{i}-y,0\right\}\big]\}\Big]$
$=$ $-\mathbb{E}_{T_{i}}\Big[\max\{0,-c_{i}^{F}+\mathbb{E}_{V_{i}\mid T_{i}}\big[(V_{i}-\sigma^{P}_{i})^{+}\big]\}\Big]+\mathbb{E}_{T_{i}}\Big[\max\{0,-c_{i}^{F}+\mathbb{E}_{V_{i}\mid T_{i}}\big[\max\left\{V_{i}-y,0\right\}\big]\}\Big].$
Similarly, the P-threshold can be interpreted as a break-even point for the value of $y$ under which P-opening is preferred to stopping, and over which stopping is preferred.
Regarding the FP-threshold $\sigma^{F/P}_{i}$, we note that the difference between the expected profits from the optimal course of actions starting from P-opening box $i$ and that from the optimal course of actions starting from F-opening box $i$ is
$\left(-c_{i}^{P}+\mathbb{E}_{T_{i}}\Big[\max\left\{y,-c_{i}^{F}+\mathbb{E}_{V_{i}\mid T_{i}}\big[\max\left\{V_{i},y\right\}\big]\right\}\Big]\right)-\left(-c_{i}^{F}+\mathbb{E}\Big[\max\left\{V_{i},y\right\}\Big]\right)$
$=-c_{i}^{P}+\mathbb{E}_{T_{i}}\Big[\max\{0,c_{i}^{F}-\mathbb{E}_{V_{i}\mid T_{i}}\big[(V_{i}-y)^{+}\big]\}\Big]\ .$ (12)
One can verify that the threshold $\sigma_{i}^{F/P}$ is a break-even point on the value of $y$ that determines whether P-opening or F-opening achieves a higher expected profit. Intuitively, in the single-box case, the definition for the set of thresholds $\sigma_{i}^{F}$, $\sigma_{i}^{F\mid t_{i}}$, $\sigma_{i}^{P}$, and $\sigma_{i}^{F/P}$ exploits the single-crossing property between pairs of action-value functions. In the more general case, the definition can be viewed as using a myopic approximation for the action-value functions that assumes that only one box may be opened.
Based on the above discussion, we construct a natural threshold-based policy for the single closed box setting. We assign the index $y$ to the action of stopping, the index $\sigma_{i}^{P}$ to the action of P-opening, and the index $\sigma^{F}_{i}$ to the action of F-opening. Our policy prescribes which action to take by comparing their corresponding indices. If the largest threshold is $y$, then we stop and choose the largest prize at hand. Otherwise, if the largest threshold is $\sigma^{F}_{i}$, then we $F$-open box $i$ and choose the largest prize at hand. Finally, if the highest threshold is $\sigma^{P}_{i}$, we compare $y$ to $\sigma^{F/P}_{i}$: if $y\leq\sigma^{F/P}_{i}$, we F-open, otherwise we P-open box $i$. We refer the reader to Figure 2 for a pictorial illustration; the policy is also summarized as Algorithm 2 in Appendix 9.
[…]
We define the largest opening threshold $\sigma_{M}(\mathcal{C},\mathcal{P})$:
$\sigma_{M}(\mathcal{C},\mathcal{P})=\max\left\{\underset{j\in\mathcal{C}}{\max}\left\{\sigma_{j}^{F},\sigma_{j}^{P}\right\},\underset{j\in\mathcal{P}}{\max}\left\{\sigma_{j}^{F\mid t_{j}}\right\}\right\}.$ (13)
We refer to the box attaining the largest threshold $\sigma_{M}(\mathcal{C},\mathcal{P})$ as the leading box.
Theorem 3.2
The threshold-based policy is optimal in the single closed box setting $({\cal C},{\cal P},y)=(\{i\},\emptyset,y)$, and it achieves the following expected profit:
[…]
In general, for every state $(\mathcal{C},\mathcal{P},y)$, it is optimal to stop if and only if
$y\geq\sigma_{M}(\mathcal{C},\mathcal{P}).$ (14)
The formal proof of the stopping rule is based on a coupling argument, which is presented in Appendix 9. We argue that the stopping decision regarding a box (i.e., never to open this box in the future) only depends on $y$. Hence, the optimal stopping rule shown for the single box setting holds for the general case.
3.2 Sufficient conditions for optimal F-opening
Having characterized the optimal stopping criterion, we now focus on states $(\mathcal{C},\mathcal{P},y)$ in which stopping is not optimal. Here, there are two possibilities: $\sigma_{M}(\mathcal{C},\mathcal{P})$ corresponds either to an F-threshold or a P-threshold. The next theorem establishes that in the former case, it is optimal to F-open the leading box. This condition holds regardless of whether the leading box is closed or partially-opened.
Theorem 3.3
There exists an optimal policy that immediately F-opens the leading box in every state $(\mathcal{C},\mathcal{P},y)$ where stopping is suboptimal and $\sigma_{M}(\mathcal{C},\mathcal{P})$ is an F-threshold.
The proof of this result appears in Appendix 9.4. The crux of it is constructing a policy that F-opens the leading box, and then showing that this policy achieves a larger expected profit than any alternative one using interchange and coupling arguments. Compared to the standard Pandora’s box problem, the main difficulty here is that P-opening introduces more courses of action to consider (e.g., P-open the leading box or alternative ones). Hence, our proof carefully distinguishes between different cases, where the leading box is P-opened and those where it is closed, and proceeds by induction. We need to quantify the value of information gathered by P-opening in order to eventually rule out this option. To this end, we upper bound this quantity using a probabilistic coupling, which exploits the dynamic programming tree locally.
We mention in passing that as part of the proof of the theorem we also show that boxes whose F-threshold is higher than their P-threshold should never be P-opened. The opposite direction does not hold, as our illustrative examples suggest.
3.3 Sufficient conditions for optimal P-opening
We consider the alternative case where $\sigma_{M}(\mathcal{C},\mathcal{P})$ is a P-threshold. Not surprisingly, in this case, there is no straightforward decision rule that prescribes the optimal action as a function of the opening thresholds. This was illustrated in Section 2.2, even in the simplest case of a single closed-box where the optimal inspection mode could well be an F-opening or P-opening action, depending on the value of $y$ (we provide another example in Appendix 7 that shows the intricacy of optimal policies). We identify a sufficient condition under which P-opening the leading box is optimal.
Definition 3.4
Given a state $(\mathcal{C},\mathcal{P},y)$ and a closed box $i\in\mathcal{C}$, we define $\sigma_{-i}=\sigma_{M}(\mathcal{C}\setminus\{i\},\mathcal{P})$. Suppose $\sigma_{i}^{P}>\sigma_{i}^{F}$. We say that box $i$ is well-classified if $y>\sigma_{i}^{F/P}$ and for every type $t_{i}\in\Gamma_{i}$, one of the following conditions holds: (1) $\sigma_{i}^{F\mid t_{i}}\geq\sigma_{-i}$ or (2) $y>\sigma_{i}^{F\mid t_{i}}$.
To unpack this notion, we remark that, when dealing with a well-classified box $i$, there is no ambiguity about the action that follows P-opening. Put simply, upon P-opening, the box can be classified as “good” or “bad.” After observing its type $t_{i}\in\Gamma_{i}$, our definition delineates two cases. In case (1), we infer from Theorem 3.3 that it is optimal to immediately F-open the same box (good case), whereas in case (2), we infer from Theorem 3.2 that box $i$ will never be F-opened (bad case). Importantly, our notions of “good” and “bad” types are preserved in later stages of the decision process, regardless of which actions are taken. As a result, for a well-classified box, it is relatively easier to quantify the value of information afforded by the P-opening.
Theorem 3.5
There exists an optimal policy that P-opens the leading box in every state $(\mathcal{C},\mathcal{P},y)$ where stopping is not optimal, $\sigma_{M}(\mathcal{C},\mathcal{P})$ is a P-threshold, and the leading box is well-classified.
The proof, presented in Appendix 9.5, is based on induction using interchange arguments. Similarly to Theorem 3.3, the analysis requires a probabilistic coupling of different cases.
Unfortunately, combining Theorems 3.2, 3.3, and 3.5 does not determine an optimal policy in every state of the DP. Unlike the optimal stopping condition, Theorems 3.3-3.5 provide sufficient but not necessary conditions. Nevertheless, in Section 5.2, we numerically solve synthetic instances and show that these theorems capture optimal decisions in most system states. Theorem 3.5 is relatively more useful in “later stages” of the decision process, where $y$ tends to be larger, and condition (2) of Definition 3.4 becomes easier to satisfy. Moreover, our theorems speed up the DP computation. When the conditions of Theorems 3.23.5 hold, there is no need to apply the DP recursion.
That said, we identify a special case in which Theorems 3.23.5 fully characterize an optimal policy. This happens when every box $i$ such that $\sigma_{i}^{P}>\sigma_{i}^{F}$, has only two types (i.e., $|\Gamma_{i}|=2$), and the FP-threshold $\sigma_{i}^{F/P}$ is negative (i.e., $\sigma_{i}^{F/P}<0$). To see this, note that whenever stopping is suboptimal, we have two cases: If the largest opening threshold is an F-threshold, it is optimal to F-open the leading box (Theorem 3.3). If the largest threshold is a P-threshold, then the leading box $i$ is well-classified. Indeed, in view of the discussion that follows Definition 3.1, we know that $\underset{t_{i}\in\Gamma_{i}}{\min}~\sigma_{i}^{F|t_{i}}\leq\sigma_{i}^{F/P}<0\leq y$ and $\underset{t_{i}\in\Gamma_{i}}{\max}~\sigma_{i}^{F|t_{i}}\geq\sigma_{i}^{P}>\sigma_{-i}$. Intuitively, this setting captures scenarios in which alternatives can be either very good or very bad. If partial opening is cheap enough, it should always be exercised. We summarize this fact in the next corollary.
Corollary 3.6
Suppose that every box $i\in[N]$ such that $\sigma_{i}^{P}>\sigma_{i}^{F}$ satisfies $|\Gamma_{i}|=2$ and $\sigma_{i}^{F/P}<0$. Then, Theorems 3.2, 3.3 and 3.5 fully determine an optimal policy.
3.4 An optimal policy for a problem with binary prizes
Next, we study a special case of PSIs whereby boxes are statistically identical and independent, contain binary Bernoulli prizes, and belong to one of two types—“good” or “bad.” We refer to this problem as PSI-B2I (Bernoulli, 2-type, identical and independent). This setting applies to situations where a decision maker is searching for a satisfactory alternative among options that are initially perceived as equivalent. Partial inspection reveals whether an alternative is promising (“good”) or has low potential (“bad”), while full inspection discloses the exact outcome. Several variants of this problem in which the decision maker cannot inspect boxes sequentially have been studied in the literature (see, e.g. Doval, 2018, and the references therein). Our setting is also different from Clarkson et al., (2020) who assume that there is exactly one prize hidden in the boxes and that opening can be done repeatedly (which does not fit our motivating applications).
In striking contrast to the problem class with binary prizes which is shown to be NP-hard (Section 4.1), the optimal policy to PSI-B2I admits a simple and intuitive structure. To formally describe it, we first remark that the equivalence of boxes implies that their opening thresholds are equal; hence, we will refer to them as $\sigma^{F}$ and $\sigma^{P}$. If $\sigma^{F}\geq\sigma^{P}$, the optimal policy is fully characterized by Theorems 3.2 and 3.3: the optimal policy F-opens closed boxes until a prize is found.
In the difficult case $\sigma^{P}>\sigma^{F}$, we denote by $n_{C},n_{G},n_{B}$ the number of closed, good and bad boxes, respectively, in the current state. Our main result in this section shows that there exists a threshold $N_{C}$ on the number of closed boxes so that when $n_{C}>N_{C}$, an optimal policy exclusively P-opens closed boxes until stopping or reaching a state in which $n_{C}\leq N_{C}$, at which point it exclusively F-opens closed boxes. We note that the threshold $N_{C}$ is fixed and independent of the state. Moreover, an optimal policy immediately F-opens good-type boxes upon discovery, while postponing bad-type boxes until all closed boxes have been at least partially opened. There are exactly $N$ such policies—corresponding to each value of $N_{C}$—that can be easily enumerated and evaluated for the purpose of optimization.
The structure of the policy above captures the case $\sigma^{P}\leq\sigma^{F}$, using the convention $N_{c}=\infty$. We also observe that the special case where $\sigma^{F}<\sigma^{P}$ and $\sigma^{F/P}\leq 0$ is subsumed by Theorem 3.5 (see Observation 1), in which case $N_{c}=0$. Therefore, to ease our analysis, we make without loss of generality the following simplifying assumption: {assumption} The boxes of PSI-B2I satisfy: $0<\sigma^{F/P}<\sigma^{F}<\sigma^{P}.$
Theorem 3.7
For a PSI-B2I instance satisfying Assumption 3.4, there exists an optimal policy $\pi$ that satisfies the following properties:
  1. 1.
    Policy $\pi$ assigns the highest priority to F-opening good-type boxes.
  2. 2.
    Policy $\pi$ either discards all bad-type boxes, or assigns the lowest priority to F-opening them.
  3. 3.
    There exists a finite threshold $N_{C}$ such that policy $\pi$ prioritizes P-opening closed boxes over F-opening them iff the number of closed boxes $n_{C}>N_{C}$.
The proof can be found in Appendix 9.6. The theorem is presented as an algorithm in Appendix 9.7.
The structure of the optimal policy suggests that the relation between the opening thresholds is the primary factor that determines whether or not P-opening should be considered. If $\sigma^{P}>\sigma^{F}$ then P-opening should be considered to identify “promising” boxes. This option is exerted whenever there are sufficiently many a priori equivalent alternatives (as captured by the threshold rule $n_{C}>N_{C}$). Intuitively, due to the uncertainty about finding good-type boxes, multiple boxes must be P-opened to obtain with high probability a satisfactory outcome. By contrast, when the total number of boxes is small, the decision-maker will most likely end up opening most of them and therefore P-opening is not beneficial.
Observe that this policy satisfies the optimal characterization given in Theorems 3.2-3.5; indeed, the proof in Appendix 9.6 directly builds on these properties. Interestingly, although PSI-B2I is a special case, much of the structure exhibited by the optimal policy for this instance extends naturally to the broader class of PSI problems. For example, it is useful to consider a class of policies that commit ex ante to fully or partially opening each closed box before execution; this notion of a committing policy proves useful in the general problem and is discussed in detail in Section 4.3. As another illustration, the instance highlights how an increase in the number of boxes makes partial opening more advantageous, a phenomenon that we further validate numerically in Section 5.4.
4 Simple Approximately Optimal Policies
We establish that computing optimal policies for PSI is NP-hard (Section 4.1). We therefore turn our attention to the development and analysis of simple approximate solutions that utilize relaxations that approximate the value function. In Section 4.2, we propose two approximation methods: Whittle’s integral and the free-info relaxation, which adapt tools from related literature. Interestingly, we show that Whittle’s integral provides a tighter upper bound, and establish the asymptotic optimality of the index policy based on the opening thresholds. In general, we exploit these problem relaxations in Section 4.3 to devise approximation algorithms, in particular achieving a (11/e)(1-1/e)-approximation ratio, through a reduction to the maximization of stochastic submodular functions (Asadpour and Nazerzadeh, 2015, Singla, 2018, Beyhaghi and Kleinberg, 2019).
4.1 Hardness of the problem
Theorem 4.1
Computing optimal policies for PSI is NP-hard.
The proof strategy involves a reduction from a family of subset sum problems. Specifically, we construct a hard instance in which every box $i$ has two types $|\Gamma_{i}|=2$: “good” and “bad.” The good type is sufficiently valuable that an optimal policy F-opens the box upon discovering it, and bad type boxes are discarded. Hence, determining the optimal policy is essentially equivalent to deciding how to partition the $N$ boxes, i.e., deciding the opening mode for each box. We carefully design this PSI instance so that the partition exactly corresponds to the solution of the subset sum problem. The proof is intricate and appears in Appendix 10.6. In particular, formalizing the preceding argument requires controlling the encoding error due to the bit precision of the input.
It is interesting to note that the hard instances we constructed for the proof of Theorem 4.1 are very close to the one subsumed by Corollary 3.6, for which we were able to provide a polynomial-time algorithm. Specifically, our family of hard instances satisfies that, every box $i\in[N]$ has $\sigma_{i}^{P}>\sigma_{i}^{F}$ and $|\Gamma_{i}|=2$. The only difference, however, is that $\sigma_{i}^{F/P}>0$, contrary to Corollary 3.6. This suggests that Corollary 3.6 is “tight” in some sense.
Finally, our reduction from subset partition has some similarities with the NP-hardness proof of Fu et al., (2023) in a related computational setting; see Remark 10.13 in Appendix 10.6 where we discuss similarities and important differences in both reduction proofs.
4.2 Relaxations and upper bounds
A common method to solve high-dimensional DPs consists in approximating the value function (see, e.g., Chapter 9 in Powell, 2007). For example, effective policies are often obtained by solving the approximate Bellman’s equation, where the true (unknown) value function is replaced by its approximate counterpart. In what follows, we construct two natural upper bounds on $J({\cal C},{\cal P},y)$ and compare them analytically.
Whittle’s integral. This is a known upper bound on the value function of MAB superprocesses (see, e.g., Brown and Smith, 2013). Consider a problem with a single non-fully-open box $i$ where the current maximum prize is $y$. Let $\omega_{i}(y)$ be the probability that the current maximum prize $y$ is eventually selected by an optimal policy. In other words, we can think of $\omega_{i}(y)$ as the probability that the effort of opening box $i$ is in vain, provided that only box $i$ is left. Given the characterization of the optimal policy for the single-box case in Theorem 3.2, it is not hard to see that, if box $i\in\mathcal{C}$,
[…]
Otherwise, if box $i\in\mathcal{P}$,
$\omega_{i}\left(y\right)=\operatorname{Pr}\left[\min\{V_{i},\sigma_{i}^{F\mid t_{i}}\}\leq y\mid T_{i}=t_{i}\right]\ .$
Consequently, Whittle’s integral corresponds to the following quantity:
$J^{W}(\mathcal{C},\mathcal{P},y)=\sigma_{M}(\mathcal{C},\mathcal{P})-\int_{y}^{\sigma_{M}(\mathcal{C},\mathcal{P})}\prod_{i\in{\cal C}\cup{\cal P}}\omega_{i}(u)du\ .$ (15)
We will establish in Lemma 4.2 that this function provides an upper bound on the $J(\mathcal{C},\mathcal{P},y)$. Conversely, when $y=\sigma_{M}(\mathcal{C},\mathcal{P})$, we straightforwardly obtain $J(\mathcal{C},\mathcal{P},y)\geq y\geq\sigma_{M}(\mathcal{C},\mathcal{P})\geq J^{W}(\mathcal{C},\mathcal{P},y)$, implying that the upper bound (15) can be tight in certain states.
Free-info relaxation.
We now consider an alternative upper-bounding method that exploits the interpretation of indices as fair strike prices of two-stage call options. To this end, we introduce indicator random variables that describe the actions chosen by any given policy $\pi\in\Pi$ from state $({\cal C},{\cal P},y)$.Recall that, in the state description, $y$ is the reward of a best F-opened box. Without loss of optimality, we restrict attention to policies that discard other $F$-opened boxes. Let $\mathbb{S}_{i}^{\pi},\widetilde{\mathbb{S}}_{i}^{\pi}$ be the binary random variables that indicate whether box $i$ is eventually selected without a partial opening and after a partial opening, respectively. (Note that we necessarily have $\mathbb{S}_{i}^{\pi}+\widetilde{\mathbb{S}}_{i}^{\pi}\leq 1$). Similarly, let $\mathbb{F}_{i}^{\pi},\widetilde{\mathbb{F}}_{i}^{\pi}$ be the binary random variables that indicate whether box $i$ is F-opened without any prior P-opening or after a P-opening, respectively. Finally, let $\mathbb{P}_{i}^{\pi}$ be the binary random variable that indicates whether box $i$ is P-opened. While these decision variables depend on the policy $\pi$, for ease of notation, we sometimes suppress the superscript $\pi$. With this notation at hand, the expected value of policy $\pi$ in state $({\cal C},{\cal P},y)$ can be written as
$J^{\pi}({\cal C},{\cal P},y)=\mathbb{E}\left[{\left.{\mathbb{S}}_{0}^{\pi}\cdot y+\sum_{i\in{\cal C}}\left(\mathbb{S}_{i}^{\pi}V_{i}-\mathbb{F}_{i}^{\pi}c_{i}^{F}-\mathbb{P}_{i}^{\pi}c_{i}^{P}\right)+\sum_{i\in{\cal C}\cup{\cal P}}\left(\widetilde{\mathbb{S}}_{i}^{\pi}V_{i}-\widetilde{\mathbb{F}}_{i}^{\pi}c_{i}^{F}\right)\right|({\cal C},{\cal P},y)}\right]\ ,$ (16)
where, by a slight abuse of notation, the conditional expectation indicates that policy $\pi$ is executed starting from state $({\cal C},{\cal P},y)$, and thus, we only accrue the costs of future opening decisions and the rewards of future selections decisions. Hence, the Pandora’s box problem with sequential inspections can be formulated as the following mathematical program:
$J({\cal C},{\cal P},y)=\max_{\pi\in\Pi}~J^{\pi}({\cal C},{\cal P},y)\ .$ (17)
The free-info relaxation considers an alternative setting, in which there is no cost of P-opening and F-opening, but instead, the rewards of the boxes are capped according to the threshold values. The notion of a feasible policy does not change in this new formulation, only the cost accounting is modified. Specifically, for every box $i\in\mathcal{C}$, we define the capped value of F-opening to be $K_{i}=\min\{V_{i},\sigma^{F}_{i}\}$ and the capped value of P-opening to be $\widetilde{K}_{i}=\min\{V_{i},\sigma^{F\mid T_{i}}_{i},\sigma^{P}_{i}\}$. Similarly, for every box $i\in{\cal P}$, we define the capped value of type $t_{i}$ as $\widetilde{K}_{i}^{t_{i}}=\min\{\sigma_{i}^{F\mid t_{i}},V_{i}\}$. Consequently, in the free-info formulation, the expected reward of policy $\pi$ in state $({\cal C},{\cal P},y)$ is
$J^{\pi,K}(\mathcal{C},\mathcal{P},y)$ $=$ $\mathbb{E}\left[\left.{\mathbb{S}}_{0}^{\pi}\cdot y+\sum_{i\in{\cal C}}\left(\mathbb{S}_{i}^{\pi}K_{i}+\widetilde{\mathbb{S}}_{i}^{\pi}\widetilde{K}_{i}\right)+\sum_{i\in{\cal P}}\widetilde{\mathbb{S}}_{i}^{\pi}\widetilde{K}^{t_{i}}_{i}\right|({\cal C},{\cal P},y)\right]\ .$ (18)
Recall that we interpret the caps as fair strike prices of call options on the box values, whose expected payoff offsets the opening costs. The formulation (18) can be viewed as an alternative cost accounting where Pandora opens boxes free of charge (there are no negative terms), and in exchange, there is a short position on a call option that exactly compensates the opening costs. However, Pandora has the advantage of being subjected to the short contract only if the box is ultimately selected; otherwise, the short contract is void. Under this optimistic assumption, Pandora essentially collects one of the capped rewards, $K_{i}$, $\widetilde{K}_{i}$, or $\widetilde{K}^{t_{i}}_{i}$, depending on the chosen opening mode. Going one step further, the free-info relaxation in state $(\mathcal{C},\mathcal{P},y)$ corresponds to the following quantity:
$J^{K}(\mathcal{C},\mathcal{P},y)=\mathbb{E}\left[\max\left\{y,\underset{i\in\mathcal{C}}{\max}\left\{\max\left\{K_{i},\widetilde{K}_{i}\right\}\right\},\underset{i\in\mathcal{P}}{\max}\left\{\widetilde{K}_{i}^{t_{i}}\right\}\right\}\right]\ .$ (19)
Similarly to Whittle’s integral, we eliminate the dependence on policy $\pi$ by choosing the box with the highest capped reward. Therefore, $J^{K}(\mathcal{C},\mathcal{P},y)$ is an upper bound on (18), which might not be attainable through a specific policy $\pi$. We are now ready to state our main result with respect to Whittle’s integral and the free-info relaxation, showing that both methods provide valid upper bounds on the value function.
Lemma 4.2
For every state $(\mathcal{C},\mathcal{P},y)$, we have
$J(\mathcal{C},\mathcal{P},y)\leq J^{W}(\mathcal{C},\mathcal{P},y)\leq J^{K}(\mathcal{C},\mathcal{P},y)\ .$ (20)
The proof appears in Appendix 10.1. These upper bounds have been independently derived in the literature for related optimization settings. To our knowledge, this result is the first to formally connect these methods and show that Whittle’s integral provides a tighter upper bound. The proof relies on results by Brown and Smith, (2013) who exhumed an unnoticed technical claim in Whittle, (1980) to derive Whittle’s integral (for the cost-discounted case), and Beyhaghi and Kleinberg, (2019) who proposed the free-info relaxation; a technical difference in our setting is that information can be acquired sequentially over multiple stages. We note in passing that Lemma 4.2 can be used to establish the asymptotic optimality of the index policy that prioritizes the highest threshold inspection. For brevity, we relegate this result to Appendix 10.2.
4.3 Near-optimal algorithms
In this section, we leverage the upper bounds of Section 4.2 to construct simple near-optimal policies. Specifically, we develop an algorithm with a worst-case approximation ratio of (11/e)(1-1/e) using a reduction to the stochastic submodular optimization problem. Our approach is an application of the framework by Singla, (2018) and Beyhaghi and Kleinberg, (2019). In this context, the main idea is to focus on policies for which the free-info upper bound $J^{\pi,K}$ of equation (18) is exactly equal to the value function $J^{\pi}$ of equation (17). It turns out there is a simple recipe for constructing such policies: (i) commit before opening any box on the selection mode of each box (i.e., whether we start by F-opening or P-opening it), and (ii) strictly follow the order of decreasing opening thresholds. That is, we partition the set of boxes into two subsets, $N]=F\cup P$, such that boxes $i\in F$ are initially $F$-opened, whereas boxes $i\in P$ are $P$-opened, if inspected. Next, in each state, we select the action with the highest associated threshold. The resulting class of policies $\pi^{F,P}$ is termed committing policies. The reader is referred to the Algorithm 1 for a complete description.
Algorithm 1 Committing policy $\pi^{F,P}$.
1:Initialize state $(\mathcal{C},\mathcal{P},y)\leftarrow([N],\emptyset,0)$;
2:while $y<\max\left\{\underset{j\in\mathcal{C}\cap F}{\max}\left\{\sigma_{j}^{F}\right\},\underset{j\in\mathcal{C}\cap P}{\max}\left\{\sigma_{j}^{P}\right\},\underset{j\in\mathcal{P}}{\max}\left\{\sigma_{j}^{F\mid t_{j}}\right\}\right\}$ do
3:  Let $\sigma_{i}=\max\left\{\underset{j\in\mathcal{C}\cap F}{\max}\left\{\sigma_{j}^{F}\right\},\underset{j\in\mathcal{C}\cap P}{\max}\left\{\sigma_{j}^{P}\right\},\underset{j\in\mathcal{P}}{\max}\left\{\sigma_{j}^{F\mid t_{j}}\right\}\right\}$;
4:  if $i\in F$ then
5:   F-open box $i$; $y\leftarrow\max\{y,v_{i}\}$; $\mathcal{C}\leftarrow\mathcal{C}\setminus\{i\}$;
6:  else if $i\in P$ then
7:   if $i\in\mathcal{C}$ then
8:     P-open box $i$; $\mathcal{C}\leftarrow\mathcal{C}\setminus\{i\};~\mathcal{P}\leftarrow\mathcal{P}\cup\{(i,t_{i})\}$;
9:   else if $i\in\mathcal{P}$ then
10:     F-open box $i$; $\mathcal{P}\leftarrow\mathcal{P}\setminus\{(i,t_{i})\};~y\leftarrow\max\{y,v_{i}\}$;
11:   end if
12:  end if
13:end while
14:Select $y$ and stop;
The next lemma shows that the expected profit of any committing policy indeed coincides with its free-info upper bound, which can be expressed in a simple form.
Lemma 4.3
Given a partition $(F,P)$ of the set of boxes $[N]$, the committing policy $\pi^{F,P}$ generates an expected profit of
$J^{\pi^{F,P}}(s_{0})=J^{\pi^{F,P},K}(s_{0})=\mathbb{E}\left[\max\left\{\underset{i\in F}{\max}\{K_{i}\},\underset{i\in P}{\max}\{\widetilde{K}_{i}\}\right\}\right].$ (21)
Intuitively, committing policies are not exposed to any cost accounting error, in comparing the opening costs to the corresponding short positions on call options. Recall that this free-info cost accounting is optimistic because Pandora is subjected to the short contract only if the box is ultimately selected. In the proof of Lemma 4.3, we identify necessary and sufficient conditions such that Pandora always selects any F-opened box (or F-opens any P-opened box) whenever the call option (or two-stage call option) is in the money. This way, Pandora meets the contractual payment of the option, which is equal to the opening costs in expectation. In particular, we show these necessary and sufficient conditions are met by committing policies. Now, this introduction of such committing policies raises the following question: how should we design the partition $(F,P)$? Can we construct committing policies that achieve strong (constant-factor) approximation ratios? Based on the expected profit characterization of Lemma 4.3, we can compare committing policy $\pi^{F,P}$’s payoff $\mathbb{E}[\max\{\underset{i\in F}{\max}\{K_{i}\},\underset{i\in P}{\max}\{\widetilde{K}_{i}\}\}]$ to the free-info upper-bound $J^{K}(s_{0})=\mathbb{E}[\underset{i\in[N]}{\max}\max\{K_{i},\widetilde{K}_{i}\}]$ (see Equation (19)). The former ex-ante chooses either $K_{i}$ or $\widetilde{K}_{i}$, before observing their realizations, depending on whether $i\in F$ or $i\in P$. By contrast, the latter chooses the maximum reward $\max\left\{K_{i},\widetilde{K}_{i}\right\}$ out of the two opening modes after seeing their realizations. This informal comparison leads to the design of simple approximation algorithms. First, note that the following upper bound on the free-info setting always holds:
$J^{K}\left(s_{0}\right)$ $=$ $\mathbb{E}\left[\max_{i\in[N]}~\max\left\{K_{i},\widetilde{K}_{i}\right\}\right]$
$\leq$ $\mathbb{E}\left[\max_{i\in[N]}\left({\mathbb{I}}[i\in F]\cdot K_{i}+{\mathbb{I}}[i\in P]\cdot\widetilde{K}_{i}\right)+\max_{i\in[N]}\left({\mathbb{I}}[i\in P]\cdot K_{i}+{\mathbb{I}}[i\in F]\cdot\widetilde{K}_{i}\right)\right]$
$=$ $J^{\pi^{F,P}}\left(s_{0}\right)+J^{\pi^{P,F}}\left(s_{0}\right)\ ,$
where the inequality holds since $\max\{a,b\}\leq\max\{c,d\}+\max\{e,f\}$ for every scalars $a\leq\max\{c,e\}$ and $b\leq\max\{d,f\}$, and the last equality proceeds from Lemma 4.3. Importantly, the above inequality implies that, for any partition of $(F,P)$ of $[N]$, either the committing policy $\pi^{F,P}$ or the one $\pi^{P,F}$ that flips the opening modes achieves a factor-$\frac{1}{2}$ approximation of the optimum. More broadly, the performance of a well-designed committing policy is related to the adaptivity gap in stochastic submodular maximization; the question can be reformulated as optimizing a certain submodular function over random entries and the goal is to bound the gap between non-adaptive policies, that select those entries before their realizations are revealed, and adaptive ones. By leveraging fundamental results in this area, we obtain the following theorem.
Theorem 4.4
There exists a committing policy that achieves at least $1-1/e\approx 0.63$ of the optimum and can be approached via a polynomial-time approximation scheme.
In Appendix 10.5, we devise a reduction from Pandora’s box problem to the problem of maximizing a stochastic submodular function under additional constraints. In this context, Asadpour and Nazerzadeh, (2015) showed, among other results, that the adaptivity gap (i.e., the performance gap between non-adaptive and adaptive policies) is lower bounded by 11/e1-1/e, and that an approximate non-adaptive policy can be efficiently constructed using a continuous-greedy algorithm.
5 Numerical Study
We conduct numerical experiments to further explore the characteristics of optimal policies and assess the effectiveness of natural candidate policies. Section 5.1 describes the process for generating a diverse set of problem instances. Section 5.2 quantifies the extent to which our theoretical characterization captures optimal decisions. Section 5.3 compares the performance of various policies inspired by theoretical analysis. Finally, in Section 5.4, we examine factors that increase the tendency of the optimal policies to P-open.
5.1 Problem instances
Our design of numerical experiments is based on the observation that the interaction between opening thresholds significantly impacts the behavior of optimal policies. We therefore create instances that consist of boxes with diverse values of opening thresholds. We begin by generating a collection of “prototypical” boxes. To create a prototypical box $i$, we sample 5 values for the support of $V_{i}$ from the uniform distribution $\mathcal{U}(0,10)$; every type $T_{i}$ can realize up to at most three types (that is, $\mid\Gamma_{i}\mid\leq 3$); for each type we uniformly sample the conditional distribution $V_{i}\mid T_{i}=t$ over the support; we sample the F-opening and P-opening costs from the uniform distributions $\mathcal{U}(0,5)$ and $\mathcal{U}(0,3)$, respectively. We then generate a candidate set of prototypical boxes out of which we select a small representative set (Figure 3). Each problem instance consists of randomized copies of the prototypical boxes, in varying compositions with $N=2$ to $N=16$ boxes. Due to running time limitations, there are 1000 instances whose size is 2 to 9 (“small instances” which can be solved to optimality using DP), and there are 300 instances whose size is 10 to 16 (“large instances” which can only be solved using heuristics). To avoid trivial cases, we only generate boxes where $c^{F}_{i}\geq c^{P}_{i}$ (which is why in Figure 3 most prototypical boxes have greater P-thresholds).
5.2 Quantifying the coverage of analytical results
The analysis of Section 3 revealed various insights about the nature of the optimal policy. Since the analysis includes sufficient but not always necessary conditions, we are interested in understanding to what extent the sufficient conditions of Theorems 1–3 describe an optimal policy. To this end, we conduct two types of experiments. First, we examine optimal solutions to the “small instances” and compute the percentage of states where the sufficient conditions of Theorems 1–3 hold (i.e., where our theorems prescribe optimal actions). We implement the DP recursion in (5) and compute coverage rates, defined as the percentages of system states where the optimal action can be determined by our sufficient conditions. The results are summarized in Table 1. The overall coverage rate is presented in the 5th column, and is divided into specific conditions in the 2nd to 4th columns. Looking at the 5th column of the table, we see that as the problem increases in size, the sufficient conditions of Theorems 1–3 capture optimal actions in all but a relatively few states. The last two columns present the recall of Theorems 2 and 3. These reflect the percentage of states in which the sufficient conditions of the theorems hold out of all states where F-opening (column 6) and P-opening (column 7) are optimal. For example, for instances of size 5 (4th row in the table), Theorem 2 captures 97% of states where F-opening is optimal, and Theorem 3 captures 77% of all states where P-opening is optimal. We observe that identifying P-opening is more challenging than identifying when F-opening is optimal, but overall, the experiment suggests that our characterization explains to a large degree the behavior of the optimal policy. Note that the recall of Theorem 1 is 100% as the theorem perfectly describes when stopping is optimal. We conduct a second experiment where we compare two DP implementations that recursively compute the optimal policy. The first policy “Naïve DP” is implemented using the DP recursion in Equation (5), which does not leverage any structure of the optimal policy. The second implementation, on the other hand, utilizes the sufficient conditions of Theorems 1–3 which determine optimal actions and do not require calling the DP recursion when the conditions are satisfied. Table 2 compares the two approaches in terms of the total number of states created, the total number of visits to system states, and the runtime. We see that the characterization of the optimal policy not only provides insights about the optimal policy, but also bears computational benefits, allowing to solve problems at significantly shorter times. That said, the “curse of dimensionality” cannot be avoided, as can be observed by the rate at which the size of the state space grows.
5.3 Policy benchmark and computational study
We evaluate and compare various policies that arise in our analysis:
  • $\pi^{\textrm{OPT}}$ – the optimal policy, computed by solving the DP naively.
  • $\pi^{F^{*},P^{*}}$ – the best committing policy, obtained through complete enumeration.
  • $\pi^{\textrm{index}}$ – an index policy defined using the opening thresholds associated with every box and action. In every state, the policy acts according to the largest index, where the thresholds $\sigma^{F}_{i}$, $\sigma^{F\mid t_{i}}_{i}$, and $\sigma^{P}_{i}$ serve as indices for the F- and P-opening actions, while $y$ is the index for selecting the best opened box. This policy is asymptotically optimal (Corollary 10.1) and satisfies the sufficient optimality conditions.
  • $\pi^{\textrm{W}}$ – a one-step lookahead policy (Bertsekas, 2019) based on approximate value functions using Whittle’s integral.
Another natural policy to consider is the one-step lookahead policy that uses $J^{K}$ to approximate the value function. However, given that $J^{W}$ provides a tighter upper bound (Lemma 4.2), we only test $\pi^{W}$ in the experiments. $\pi^{STP}$ – a policy that determines its next action based on the maximal myopic improvement (using a single “test,” hence Single Test Policy) in comparison to stopping. Computing the improvement from each non-open box becomes tractable by ignoring other non-open boxes, and follows the optimal policy for the single box problem (see Figure 2). To evaluate performance, we measure the value function and runtime of each problem instance under different policies, starting from the initial state. As the expected profit can vary significantly among different instances, we normalize the expected profit of each policy relative to the optimal value obtained by policy $\pi^{\textrm{OPT}}$. For large instances where the optimal policy cannot be computed, we use the best-performing policy to normalize the expected profit. Thus, the performance of each policy is reported as a number between 0 and 1, where larger values indicate better performance. For each policy and a set of instances, we report on performance using the average across instances and the worst instance tested. Average performance. All heuristics perform quite well in terms of their average performance, often exceeding 90% of OPT. However, $\pi^{\textrm{index}}$ and $\pi^{F^{*},P^{*}}$ are superior, with average performances that are close to 100%. Worst-case performance. We observe that $\pi^{\textrm{W}}$ can perform poorly, at times nearing zero profit. This suggests that Whittle’s integral approach may not be a suitable approach to deriving tight bounds for our problem, and perhaps to other Pandora’s box problems as well. We also see that $\pi^{\textrm{index}}$ is superior to $\pi^{STP}$ and that its relative performance improves as the number of boxes increases. We observe that $\pi^{\textrm{index}}$ approaches optimality as $N$ increases, which is aligned with its asymptotic optimality. Outperforming all others is policy $\pi^{F^{*},P^{*}}$, which achieves the best worst-case performance, with a loss of at most 7% for instances of size 2. This suggests that index policies based on our opening thresholds are effective for solving the problem. We note, however, that this policy is not optimal, and finding the optimal committing policy is computationally challenging (an approximation algorithm is provided in Asadpour and Nazerzadeh, 2015). Runtime. Table 5 in Appendix 11 summarizes the runtimes. As expected, the time for solving instances to optimality grows exponentially fast, making it challenging to solve instances with more than 9 boxes. We also see that applying policy $\pi^{\textrm{W}}$ becomes more challenging as the number of boxes increases, due to the calculation of Whittle’s integral. Similarly, the exhaustive enumeration carried out by policy $\pi^{F^{*},P^{*}}$ could result in long runtimes. We note that the runtime also depends on the size of the support of the distribution (which we assumed to be 5) and the number of box types (which we assumed to be at most 3). Note that these runtimes include the entire calculation of the value function and that simply applying these policies is very fast. Optimality. Table 4 shows the percentage of instances for which different policies attain the optimal expected objective value. In contrast to the near-optimal performance displayed in Table 3, policy $\pi^{\textrm{W}}$ is almost never optimal. Additionally, even policies $\pi^{F,P}$ and $\pi^{\textrm{index}}$, which perform better, in many cases do not achieve the optimal objective value. We notice that overall, the performance of the latter two index policies improves with the number of boxes, which aligns with the asymptotic optimality result (Corollary 10.1).
5.4 When is P-opening worthwhile?
We next conduct experiments that shed light on the structure of optimal policies and further characterize when P-opening is worthwhile. To this end, we define the notion of P-ratio as the average percentage of P-openings conducted by the optimal policy of all opening actions. A higher P-ratio implies that the optimal policy applies P-opening more frequently. We compute the P-ratio for a set of problem instances and plot it against various problem characteristics. For brevity, we describe the main results below and postpone the complete details of the experiment and the figures to Appendix 12. One of our main findings is that greater homogeneity of boxes in terms of their opening thresholds (i.e., boxes $i$ and $j$ are “homogeneous” when the tuples $(\sigma^{F}_{i},\sigma^{P}_{i})$ and $(\sigma^{F}_{j},\sigma^{P}_{j})$ are close) is correlated with more frequent partial opening. This aligns with intuition, since when boxes are very different, their relative value can be determined without partial inspection; hence, F-opening can simply be conducted as a necessity prior to selection. Additionally, in line with our theoretical results, we observe an increased tendency to partially open when there is a higher proportion of boxes whose P-thresholds exceed their corresponding F-thresholds. This finding underscores the role of opening thresholds in capturing the value of corresponding inspection modes. Finally, in the case of identical boxes (homogeneity is maximal), the tendency to partially open increases in the number of boxes. This suggests that when the pool of alternatives is larger, partial opening becomes more significant—assuming partial opening is worthwhile to begin with.
6 Conclusions and Future Directions
[…]